# MessIRve: A Large-Scale Spanish Information Retrieval Dataset

### v1

@@ 4.1 Analysis of topics
Figure 1: Topics in Spanish IR datasets. Topics are sorted by overall frequency from top to bottom, and are labeled with an ID and the four most representative words according to their c-TF-IDF values (see Grootendorst, 2022). Numbers in the heatmap represent the % of queries in a given dataset that belong to a given topic. The color scale is shared across datasets. For the sake of clarity, the names of topics were translated from Spanish to English by the authors.
To more thoroughly analyze the topics covered by the queries of the different datasets, we employed BERTopic with PCA of 30 dimensions of multilingual-e5-large embeddings, and KMeans clustering with 30 clusters (Grootendorst, 2022; Wang et al., 2024). We assume just one topic per query, and focus only on the queries of the general-domain datasets: MIRACL, SQAC, mMARCO, and MessIRve full. For every dataset we use the complete set of available queries.
The clusters were labeled with the four most representative words according to their c-TF-IDF values (Grootendorst, 2022). To provide a summary of a particular topic in a given dataset, representative queries were found by sampling a subset of queries from the dataset belonging to the topic and calculating which ones were closest to the topic’s c-TF-IDF representations based on cosine similarity. Representative words of a topic in a dataset were also found by considering the highest c-TF-IDF scores of the words from the dataset.
@@
MIRACL and SQAC show a higher frequency of topics related to war and politics (topic 19), religion and the Bible (topic 24), and inventions or discoveries (topic 25). This can be attributed to their dataset construction methodology, which starts from Wikipedia passages rather than user queries. Compared to other datasets, MessIRve shows a prominence of queries related to users experiencing health-related pains (topic 2), while mMARCO features topics related to household economy (topics 5 and 6) and computers and technology (topic 8).
It is interesting to note that topic 23, relatively prevalent in mMARCO but not in MessIRve, focuses on US demographics. This topic features representative documents such as “dallas nc se encuentra en qué condado” (Dallas, NC is located in what county), “en qué condado se encuentra grove city ohio” (In what county is Grove City, Ohio located), and “cual es el condado de washington tx” (Which is the county of Washington, TX).
This contrasts with the queries from topic 4, which are relevant in MessIRve, SQAC, and MIRACL, and focus on demographic issues. Representative queries from this topic in MessIRve include “en donde queda ciudad del este” (where is Ciudad del Este), “donde queda logroño” (where is Logroño), and “por donde queda montañita” (where is Montañita). These queries are likely to be more representative of the information needs of Spanish speakers than those in topic 23 of mMARCO, which probably come from English-speaking users.
Similarly, the queries in topic 0 of mMARCO and MessIRve highlight differences in popular entities within Spanish-speaking countries. For example, a representative mMARCO query is “en qué equipo juega michael sam” (what team does Michael Sam play for), referring to a US NFL player, while one from MessIRve is “en cuál equipo juega james rodriguez” (which team does James Rodriguez play for), referring to a Colombian soccer player. This behaviour is expected because MessIRve queries are obtained from Google’s autocomplete API by country domains, while mMARCO queries are derived from English-speaking users’ queries and translated into Spanish.
Finally, the distribution of topics in MessIRve and mMARCO is more uniform than in MIRACL and SQAC, which have a few topics that are much more frequent than others. This is likely a result of the larger size of MessIRve and mMARCO, allowing for a wider variety of topics.
However, MIRACL and SQAC contain well-formed questions, while MessIRve includes more colloquial and informal questions, which better represent the queries that users might issue to a search engine. This feature is also present in mMARCO, but affected by translation quality. See Table 5 in Appendix A for further details on representative words and documents for the most frequent topics.
We stress that the topic analysis of MessIRve’s queries is not intended to provide a faithful representation of the real distribution of popular topics for the Spanish-speaking community. Instead, it is just a descriptive analysis of the dataset, showcasing the kinds of topics included and their relative importance as compared to other datasets.
@@ 5 Experiments
We evaluate various IR models on MessIRve’s test sets in a zero-shot manner. In section 5.1 we describe the experimental setup, and in section 5.2 we present the results.
5.1 Experimental setup
We consider the following models for information retrieval in Spanish:
BM25
(Robertson et al., 2009). BM25 is a popular lexical IR model based on the token overlap between query and documents. We index the corpus and run retrieval using Pyserini with default parameters and Spanish analyzer (Lin et al., 2021). Despite its simplicity, BM25 is a strong baseline for IR (Thakur et al., 2021).
MIRACL-mdpr-es
(Zhang et al., 2023). This is a bi-encoder initalized from multilingual Dense Passage Retriever (mDPR, Zhang et al., 2021b), which is first pre-fine-tuned using the training set of MS MARCO and further fine-tuned on the MIRACL Spanish training set. We use the mdpr-tied-pft-msmarco-ft-miracl-es checkpoint. [^5] https://huggingface.co/castorini/mdpr-tied-pft-msmarco-ft-miracl-es As far as we know, this is the only well-documented, Spanish-specific retriever available.
E5-large
(Wang et al., 2024). This bi-encoder is initialized from the multilingual model XLM-R (Conneau et al., 2020) and then trained with weak supervision in a mixture of multilingual text pairs, followed by supervised fine-tuning in annotated data which is mostly in English but includes some multilingual data e.g. MIRACL. We use the multilingual-e5-large checkpoint. [^6] https://huggingface.co/intfloat/multilingual-e5-large
OpenAI-large
. Although the architecture and training data of this model are not publicly disclosed, previous models (Neelakantan et al., 2022) consist of a single encoder that maps documents and query to embeddings, and is trained with contrastive learning across several supervised datasets. The model we use, text-embedding-3-large, is accessible via an API. [^7] https://platform.openai.com/docs/guides/embeddings
In all cases, we append the title of the Wikipedia article to the document text before retrieval to provide additional context.
We use two standard metrics to measure retrieval performance:
@@ 5.1 Experimental setup
All evaluations are conducted in a zero-shot setting, where the models are not fine-tuned on any training set. We leave this for future work.
5.2 Results
Figure 2: Baseline results on MessIRve test sets.
The OpenAI-large and E5-large models outperform smaller models (MIRACL-mdpr-es) and the lexical model (BM25) across all dataset partitions (see Figure 2). In terms of nDCG@10, OpenAI-large achieves an average score in the full set of .433 compared to E5-large’s .430, indicating similar ranking effectiveness. However, for Recall@100, OpenAI-large consistently outperforms E5-large, with a recall of .899 versus .847.
BM25 consistently performs the worst across all subsets (nDCG@10=.167 and Recall@100=.530 in the full set), highlighting the limitations of lexical retrieval methods compared to dense models.
Although MIRACL-mdpr-es is specifically fine-tuned for Spanish, it still lags behind the larger models, with nDCG@10 and Recall@100 scores of .259 and .626, respectively, suggesting that model size and fine-tuning strategy might play a more critical role than language-specific tuning. For example, while MIRACL-mdpr-es has 178M parameters and is pre-fine-tuned with MS MARCO and then fine-tuned with Spanish MIRACL, E5-large has 560M parameters, is pre-trained in a mixture of multilingual text pairs, and then fine-tuned in at least 10 supervised datasets, including MS MARCO and multilingual data like MIRACL itself.
Finally, the three dense models follow a similar performance pattern, with the highest scores in the bo (Bolivia) and gt (Guatemala) subsets and the lowest in ar (Argentina) and mx (Mexico).
@@ 6 Discussion and conclusions
We performed a quality assessment to ensure the reliability of the dataset, and we provided a detailed comparison with existing Spanish IR datasets, highlighting the dataset’s strengths and limitations. We also provide baseline results of the performance of various IR models on the dataset.
It must be noted that the dataset was collected using an IR system in itself, Google’s “featured snippets”. Although there’s no certainty about how Google’s IR system works, it is likely that it considers many other factors besides the text of the query and the documents, such as inbound links, user behavior, the structure of the web, etc., implying significant computational resources. This probably contributes to the high quality of the dataset, and so makes it a valuable resource for training and evaluating open-source IR models. However, it also introduces an evaluation bias towards Google’s own IR system, as the dataset may not be used to identify IR systems that are better than Google’s own system.
In this work, we have focused on zero-shot baselines. Future work includes fine-tuning the models on the training set of MessIRve, evaluating models on other Spanish IR datasets, and creating a unified Spanish IR benchmark with all datasets.
@@
The quality assessment of the dataset was performed by four of the authors, which may introduce bias. This was done due to the lack of resources to hire external raters.
The dataset should not be used for commercial purposes.
@@ Acknowledgments
This work used computational resources from CCAD – Universidad Nacional de Córdoba (https://ccad.unc.edu.ar/), which are part of SNCAD – MinCyT, República Argentina, and from NodaIA San Francisco (Ministry of Science and Technology of the Province of Córdoba, Argentina).
@@
We thank OpenAI for their financial support to run experiments with their models. We also thank authors of GooAQ (Khashabi et al., 2021) and Pyserini (Lin et al., 2021) for publicly sharing their code, which significantly contributed to our research.
Appendix A Supporting information

### v2

@@ 4.2 Analysis of topics
Figure 2: Topics in Spanish IR datasets. Topics are sorted by overall frequency from top to bottom, and are labeled with an ID and the four most representative words according to their c-TF-IDF values (see grootendorst2022bertopic). Numbers in the heatmap represent the % of queries in a dataset that belong to a topic. The color scale is shared across datasets. For the sake of clarity, words were translated from Spanish to English by the authors.
To more thoroughly analyze the topics covered by the queries of the different datasets, we employed BERTopic with PCA of 30 dimensions of multilingual-e5-large embeddings, and KMeans clustering with 30 clusters (grootendorst2022bertopic; wang2024multilingual). We assume just one topic per query, and focus only on the queries of the general-domain datasets: MIRACL, SQAC, mMARCO, and MessIRve full. For every dataset we use the complete set of available queries.
The clusters were labeled with the four most representative words according to their c-TF-IDF values (grootendorst2022bertopic). To provide a summary of a particular topic in a given dataset, representative queries were found by sampling a subset of queries from the dataset belonging to the topic and calculating which ones were closest to the topic’s c-TF-IDF representations based on cosine similarity. Representative words of a topic in a dataset were also found by considering the highest c-TF-IDF scores of the words from the dataset.
@@
MIRACL and SQAC have more topics on war and politics (topic 19), religion (topic 24), and inventions or discoveries (topic 25), likely due to their Wikipedia-based construction. Relative to other datasets, MessIRve focuses on health-related pain (topic 2), while mMARCO features household economy (topics 5 and 6) and technology (topic 8).
It is interesting that topic 23, prevalent in mMARCO but not MessIRve, concerns US demographics, with representative queries like “dallas nc se encuentra en qué condado” (Dallas, NC is located in what county) or “en qué condado se encuentra grove city ohio” (In what county is Grove City, Ohio located). In contrast, topic 4, relevant in MessIRve, SQAC, and MIRACL, focuses on broader demographics, with queries such as “donde queda logroño” (where is Logroño), or “por donde queda montañita” (where is Montañita), reflecting Spanish-speaking users’ needs more than topic 23 of mMARCO, whose queries probably come from English-speaking users.
Similarly, topic 0 in MessIRve and mMARCO highlights popularity differences in Spanish-speaking regions. mMARCO queries include “en qué equipo juega michael sam” (what team does Michael Sam play for), referring to a US NFL player, whereas MessIRve has “en cuál equipo juega james rodriguez” (which team does James Rodriguez play for), referencing a Colombian soccer player. This aligns with MessIRve’s extraction of queries by Spanish-speaking country, unlike mMARCO’s translated English queries.
MessIRve and mMARCO have a more balanced topic distribution than MIRACL and SQAC, where certain topics dominate in frequency. This likely results from their larger size, allowing for greater topic diversity. However, MIRACL and SQAC contain well-formed questions, whereas MessIRve has more colloquial queries, better representing real user searches. mMARCO also shares this trait but is influenced by translation quality. Table 5 in Appendix C provides representative words and queries for the most frequent topics.
We stress that MessIRve’s topic analysis is not a faithful representation of Spanish-speaking users’ real interests, but rather a descriptive overview of the dataset composition compared to others.
@@ 5 Experiments
As a baseline, we evaluate IR models on MessIRve’s and other IR datasets’ test sets. We consider the following models for evaluation:
  • BM25 (robertson1994okapi), a popular lexical IR model based on the token overlap between query and documents. Despite its simplicity, BM25 has proven to be a strong baseline (thakur2021beir).
  • MIRACL-mdpr-es (zhang2023miracl). This is a bi-encoder initalized from multilingual Dense Passage Retriever (mDPR, zhang-etal-2021-mr), which was first pre-fine-tuned using the training set of MSMARCO and further fine-tuned on the MIRACL Spanish training set. We use the mdpr-tied-pft-msmarco-ft-miracl-es checkpoint.666huggingface.co/castorini/mdpr-tied-pft-msmarco-ft-miracl-es As far as we know, this is the only well-documented, Spanish-specific retriever available.
  • E5-large (wang2024multilingual). This bi-encoder is initialized from the multilingual model XLM-R (conneau-etal-2020-unsupervised) and then trained with weak supervision in a mixture of multilingual text pairs, followed by supervised fine-tuning in annotated data which is mostly in English but includes some multilingual data e.g. MIRACL. We use the multilingual-e5-large checkpoint.777huggingface.co/intfloat/multilingual-e5-large
  • OpenAI-large. Although the architecture and training data are not publicly disclosed, previous OpenAI models (neelakantan2022text) consist of a bi-encoder that maps documents and query to embeddings, and is trained with contrastive learning on several supervised datasets. We use text-embedding-3-large via an API.888platform.openai.com/docs/guides/embeddings
  • E5-large-ft-messirve. We fine-tuned an E5-large model on the MessIRve full training set, retrieving hard negatives with BM25 and following the same approach as wang2024multilingual. Refer to Appendix A for more details on the training.
In all cases, we append the title of the Wikipedia article to the document text before retrieval. Except for E5-large-ft-messirve, all evaluations are conducted in a zero-shot manner.
We use two standard metrics for evaluation (thakur2021beir; zhang2023miracl):
@@
Figure 3: Baseline results on MessIRve test sets. Metrics include nDCG@10 and Recall@100. All dense models consistently outperform BM25. Fine-tuned E5-large achieves the highest nDCG@10 across all subsets, while zero-shot OpenAI-large leads in Recall@100.
Figure 4: Baseline results on Spanish IR datasets. We consider only open-source models due to the cost of evaluating OpenAI-large. We use the test queries of each dataset except MIRACL, for which we use the dev set, since the test set is not publicly available.
Figure 3 shows the results of the models on the MessIRve test split in terms of nDCG@10 and Recall@100, segmented by country plus the full and none sets. The OpenAI-large and E5-large models outperform smaller models (MIRACL-mdpr-es) and the lexical model (BM25) across all dataset partitions. In terms of nDCG@10, OpenAI-large achieves an average score in the full set of .476.476 compared to E5-large’s .463.463, indicating a slightly higher zero-shot ranking effectiveness. For Recall@100, OpenAI-large consistently outperforms E5-large accross the dataset partitions, with a recall of .916.916 versus .865.865 in the full set.
The fine-tuned version of E5-large on MessIRve surpasses all other zero-shot models in terms of nDCG@10 across all subsets, achieving a score of .491.491 in the full set. This means that, even if the queries in the test set refer to articles that are not in the training set, the fine-tuned model is able to learn a better ranking function. While the fine-tuning improves the zero-shot recall of E5-large to .887.887, it does not surpass the recall of OpenAI-large.
BM25 consistently performs the worst across all subsets (nDCG@10=.179.179 and Recall@100=.558.558 in the full set), highlighting the limitations of lexical methods compared to dense models, particularly in queries written in colloquial language.
Although MIRACL-mdpr-es is specifically fine-tuned for Spanish, it lags behind the larger models, with nDCG@10 and Recall@100 scores of .284 and .658, respectively, suggesting that model size and fine-tuning strategy might play a more critical role than language-specific tuning. For example, while MIRACL-mdpr-es has 178M parameters and is pre-fine-tuned with MSMARCO and then fine-tuned with Spanish MIRACL, E5-large has 560M parameters, is pre-trained in a mixture of multilingual text pairs, and then fine-tuned in at least 10 supervised datasets, including MSMARCO and multilingual data like MIRACL itself.
All dense models follow a similar performance pattern, with the highest scores in the bo (Bolivia) and gt (Guatemala) subsets and the lowest in ar (Argentina) and mx (Mexico).
We also compare the methods across the datasets described in Section 2 (Figure 4). The model rankings vary depending on the dataset, based on both Recall and NDCG. Overall, E5-large performs relatively well across all datasets. In SQAC, Multi-EuP, and PRES, BM25 also shows strong results, suggesting that lexical overlap is important in these datasets, unlike the others. The fine-tuned models, MIRACL-mdpr-es and E5-large-ft-messirve, perform best only in their respective datasets but struggle in others. This suggests that each dataset is quite different from the others. It also indicates that individual training datasets alone may not be enough for achieving generalization to other domains through fine-tuning, but they could be useful when combined with other training datasets.
Finally, to better understand the dataset’s limitations, we conducted a manual error analysis on 30 queries where the top-performing retriever, OpenAI-large, failed to find a relevant passage within the top-20 results. For each query, we reviewed the annotated relevant passages alongside the top 5 retrieved passages. We found that 56.7% (17/30) of the cases included false negatives due to non-exhaustive annotations, 23.3% (7/30) came from ambiguous queries, and 13.3% (4/30) had documents erroneously marked as relevant. This means that errors can come from dataset issues, as well as from models’ capabilities.
@@ 6 Discussion and conclusions
We performed a quality assessment to ensure the reliability of the dataset, and we provided a detailed comparison with existing Spanish IR datasets, highlighting the dataset’s strengths and limitations. We also provide baseline results of the performance of various IR models on the datasets.
It should be noted that the dataset was collected using an IR system, Google’s “featured snippets”. Although there’s no certainty about how it works, it’s likely that it considers many other factors besides the text of the query and the documents, such as inbound links, user behavior, the structure of the web, etc., implying significant computational resources. This probably contributes to the high quality of the dataset, making it a valuable resource for training and evaluating open-source IR models.
However, relying on Google’s system introduces potential biases. Our dataset may reflect Google’s ranking priorities and content preferences, and is unsuitable for identifying IR systems that outperform Google’s own technology. Like all IR datasets, ours has limitations in terms of scope, collection methods and underlying sources. To mitigate such biases, we advocate evaluating IR models using a diverse set of datasets rather than just one, as demonstrated by benchmarks such as BEIR (thakur2021beir). By contributing to dataset diversity and complementing existing resources, MessIRve supports the broader goal of building a comprehensive Spanish-language IR benchmark.
We have focused on zero-shot baselines and fine-tuning on the full set. Future work includes fine-tuning on country-specific subsets, and creating a unified Spanish IR benchmark with all datasets.
@@
While we made an effort to ensure that the dataset is free from offensive content, we did not conduct a systematic analysis. Instead, we checked for the absence of popular offensive terms in Spanish, which may have overlooked some inappropriate content.
Despite our efforts to include different Spanish-speaking regions, certain dialects or linguistic variations may still be underrepresented in the queries of the dataset. Reliance on Wikipedia as a source of documents may also introduce bias, as Wikipedia content is far from comprehensive in terms of topics and viewpoints, and may exclude topics of importance to some communities.
To examine this issue further, we analyzed Wikipedia edit statistics, specifically the average distribution of monthly active editors from 2018 to 2024. The countries contributing more than 1% include es (28.1%), ar (13.1%), mx (12.4%), cl (8.0%), co (7.4%), pe (6.3%), us (2.4%), uy (2.1%), ec (2.1%), cr (1.2%), bo (1.1%).999dumps.wikimedia.org/other/geoeditors/readme.html This distribution suggests that while contributions are skewed towards Spain, it still reflects representation from various Spanish-speaking countries, but still excluding some.
Although our dataset represents an improvement over what is currently available in Spanish, IR systems trained on this dataset could perpetuate existing biases from the queries, corpus selection, and relevance annotations of Google’s “featured snippets”. For example, retrieval results could prioritize topics or perspectives from well-represented communities, while underrepresenting minority viewpoints.
We did not assess the statistical significance of our results or compute confidence intervals. However, given the dataset’s scale, we believe that the performance estimates provided are reliable.
Finally, the dataset released should not be used for commercial purposes.
@@ Acknowledgments
This work used computational resources from CCAD – Universidad Nacional de Córdoba (https://ccad.unc.edu.ar/), which are part of SNCAD – MinCyT, República Argentina, and from NodoIA San Francisco (Ministry of Science and Technology of the Province of Córdoba, Argentina).
@@
We thank: OpenAI for their financial support to run experiments with their models; authors of GooAQ (khashabi-etal-2021-gooaq-open) and Pyserini (Lin_etal_SIGIR2021_Pyserini) for publicly sharing their code, which significantly contributed to our research; and Chris Petrillo of Wikipedia for generously providing Wikipedia edit statistics.
Appendix A Inference and fine-tuning
To evaluate the performance of BM25, we indexed the corpus and ran retrieval using Pyserini with default parameters and Spanish analyzer (Lin_etal_SIGIR2021_Pyserini).
Inference and training of dense retrieval models was performed with the transformers library version 4.35.2 (wolf2020transformers) on 2 NVIDIA A30 GPUs with 24GB of memory each. We used the deepspeed library (rasley2020deepspeed), which allows for efficient training of large models.
To fine-tune the E5-large model on MessIRve, we followed wang2024multilingual and wang2024text, and used the following hyperparameters: we train for 2 epochs using a batch size of 128, a learning rate of $1e^{-5}$ with linear decay and the first 400 steps for warmup, and we used 7 hard negatives extracted with BM25 for each positive example.
To reduce GPU memory usage, we used mixed precision training, gradient checkpointing, and truncated queries and documents to a maximum length of 64 and 256 tokens, respectively. We also keep position embeddings frozen during fine-tuning.
It took around 15 hours to fine-tune the model on the MessIRve full training set. No hyperparameter search was performed.
Appendix B Quality assessment
To provide additional support for the quality assessment of the dataset, four of the authors of the paper, native Spanish speakers from Argentina, evaluated a larger number of samples, trying to be as objective as possible and working independently of each other. Each of them rated 200 query-document pairs, with 50 pairs in common across all raters, for a total of 650 unique pairs.
93.5% of the queries were rated as correct, 90.1% were rated as unambiguous, and 90.2% of the documents were considered relevant to their queries, using the majority vote for overlapping queries. These results validate the results obtained with the hired independent raters. We also measured the inter-rater agreement by computing the % of the 50 common queries on which all annotators agree. We obtained 78% for query correctness, 74% for query unambiguity, and 78% for document relevance. These results further support the reliability of the quality assessment.
To better understand inter-rater agreement, we show the distribution of positive ratings across queries for each aspect, based on assessments by Prolific raters:
  • Correct.: 3 raters: 80%, 2: 13 %, 1: 7%, 0: 0%
  • Non-ambiguity: 3: 64%, 2: 24%, 1: 11%, 0: 1%
  • Relevance: 3: 83%, 2: 14%, 1: 3%, 0: 0%
This shows that full agreement on non-relevance is rare: all raters never simultaneously assign a negative rating. This suggests that when a passage is rated as non-relevant by some rater, it might be due to ambiguity rather than clear irrelevance.
Additionally, we see that because samples appear to be generally good (positives are the majority class), there are few samples left to evaluate what the agreement might be in the negative class. This imbalance can inflate simple % agreement: when the marginal probability of assigning a positive label is high, % agreement can be high even if ratings are independent.
To address this, we computed Fleiss’ kappa ($\kappa$) in the Prolific-rated samples: we obtained 0.186 for correctness, 0.146 for non-ambiguity, and 0.089 for relevance, indicating slight agreement. However, given the high marginal probability and the sample size, the variance of $\kappa$ is high, requiring a larger sample for reliable measurement. Intuitively, when most samples belong to one class, fewer are available to assess actual agreement.
In our analysis of instances annotated as non-relevant by GPT-4o and Claude-3.5-Sonnet, we found some queries that were inherently ambiguous, making it difficult to determine whether the retrieved document was relevant. Examples include:
  • “de cuántos grados fue el temblor de hoy en Colombia” (how many degrees was today’s earthquake in Colombia). The document retrieved discusses the Urrao Earthquake of 2021 which we know is not relevant because of the time frame of the query.
  • “quién se queda con el perro” (who gets the dog). The passage retrieved is about “¿Con quién se queda el perro?”, the title of a music album, rather than addressing a legal or situational question regarding pet ownership.
  • “qué hace crema para peinar” (what hair styling cream do). This phrase is grammatically incorrect and lacks specificity, making it difficult to determine the correct interpretation.
  • “por qué las articulaciones suenan” (why do joints/articulations make a sound). The passage retrieved discusses phonetics, while the query is more likely referring to the physiological phenomenon of joint cracking.
Other queries seem to be too general to be accurately matched with a specific document, for example “cuáles son derechos económicos” (which are economic rights); the query is broad and may refer to various aspects of economic rights, making it challenging to determine whether a given document sufficiently addresses the intended scope.
We observed cases of clear mismatch where the retrieved passage was clearly non-relevant to the query. Examples include:
  • “por qué es rosa la sal del Himalaya” (why is Himalayan salt pink). The passage retrieved discusses black salt and explicitly states that it should not be confused with pink Himalayan salt.
  • “en qué isla se filmó La Laguna Azul” (on which island was The Blue Lagoon filmed). The passage retrieved describes Blue Lagoon Island in the Bahamas but does not explicitly confirm whether this was the filming location of “La Laguna Azul”.
Finally, we identified instances where the negative annotation assigned by the LLMs might be debatable. For example, for the query “qué hizo John Wesley” (what did John Wesley do), the passage states that John Wesley was famous for organizing the 1869 expedition through the Green and Colorado Rivers in Utah. While this passage does not provide a comprehensive biography, it still presents a relevant historical action associated with John Wesley, making its exclusion questionable.
These findings highlight potential areas for refinement in future dataset evaluations.
Appendix C Queries
Figure 5: Entropy and number of unique prefixes of length 1 to 6 in each dataset. The entropy measures the diversity of the prefix distribution, with higher values indicating more variability. The number of unique prefixes indicates how many different prefixes of a given length exist in the dataset.